# Differential Contributions of Machine Learning and Statistical Analysis to Language and Cognitive Sciences

### v1

@@ Abstract
Data-driven approaches have revolutionized scientific research. Machine learning and statistical analysis are commonly utilized in this type of research. Despite their widespread use, these methodologies differ significantly in their techniques and objectives. Few studies have utilized a consistent dataset to demonstrate these differences within the social sciences, particularly in language and cognitive sciences. This study leverages the Buckeye Speech Corpus to illustrate how both machine learning and statistical analysis are applied in data-driven research to obtain distinct insights. This study significantly enhances our understanding of the diverse approaches employed in data-driven strategies.
Keywords: data-driven approaches, features, objectives, mixture,
@@ 1 Introduction
Advancements in technology have revolutionized the ability of scientists to gather and analyze data on an unprecedented scale. This surge in data availability has led to the necessity for more sophisticated analytical techniques and computational tools, fundamentally changing the way research is conducted and facilitating new discoveries across various domains (Langley, , 1981; Han et al., , 1993; Kitchin, , 2014; Montáns et al., , 2019; Jack et al., , 2018; Núñez et al., , 2019).
@@
Data technologies serve as powerful illuminators, revealing hidden patterns and insights within vast datasets, much like a flashlight cutting through darkness. For example, machine learning has become indispensable in scientific inquiry, allowing researchers to decipher complex patterns, enhance their understanding of intricate phenomena, and make precise predictions. This technology is applied across diverse fields including medicine, astronomy, genomics, social sciences, and environmental studies, enabling researchers to swiftly extract significant insights from large volumes of data, much faster than traditional methods would permit (Mjolsness and DeCoste, , 2001; Jordan and Mitchell, , 2015; LeCun et al., , 2015; Webb et al., , 2018). Such acceleration not only deepens the exploration of research questions but also unveils knowledge that was once concealed or unreachable. Furthermore, statistical methods play a critical role in the modern landscape of scientific research. They are integral in every phase of a study, from planning and design to data collection, analysis, and the interpretation and reporting of results (Carleo et al., , 2019; Núñez et al., , 2019; Butler et al., , 2018; Grimmer et al., , 2021). While machine learning helps in identifying patterns that may elude human detection in massive datasets, statistical techniques continue to enrich scientific research by providing robust frameworks for making inferences and validating findings (Fisher, , 1955; Box, , 1976; Nelder, , 1986; Nosek et al., , 2012). Together, these tools are transforming scientific paradigms, propelling forward our capacity to understand and manipulate the natural world.
@@
Data-driven research has made significant strides across various fields, bringing to the forefront disciplines such as data science, statistics, machine learning, and deep learning (Solomatine and Ostfeld, , 2008; Wolf, , 2010; Miller and Goodchild, , 2015; Zhang et al., , 2011). Among these, machine learning and statistics form the foundational core. While often conflated due to their shared capability to process data, machine learning and statistics frequently overlap, creating ambiguity around their distinct roles and unique contributions. Moreover, both machine learning and statistics take advantage of the shared knowledge of probability theory, linear algebra etc. Many researchers have found it challenging to distinguish clearly between these two areas. Although both fields are fundamentally concerned with extracting knowledge from data, they differ significantly in their approaches and objectives.
The rapid advancement of statistics in the early 20th century significantly enhanced scientists’ ability to quantitatively evaluate hypotheses such as “does treatment X affect outcome Y?”. These inquiries often required definitive, binary answers, leading to the development of statistical tools focused primarily on validating binary propositions through hypothesis testing and generating confidence intervals to estimate the magnitude and uncertainty of observed effects. Contrastingly, machine learning seemingly emerged from the field of engineering, driven by the ambition to enhance machine functionality. Initially, machine learning aimed to augment machine capabilities, with a secondary focus on comprehending intelligence itself. In this discipline, verifying real-world truths ranked lower in priority, considered relevant mostly as components of intelligent behaviors, a point still open to debate.
@@
While both statistics and machine learning are interconnected, they hold distinct identities, as evidenced by various studies. Gaining a clear understanding of these differences is crucial for mastering both fields and effectively applying data-driven strategies. Contemporary research methodologies frequently incorporate advanced techniques such as machine learning and statistical analysis, often leveraging both to harness the strengths of each approach. Despite the prevalent use of these methods, there remains a notable gap in studies that systematically apply both techniques to identical datasets to highlight potential differences in outcomes. This comparative approach is rarely employed. However, it could provide valuable insights into how each method processes and interprets the same dataset differently. Moreover, the specific contributions of these techniques to various academic fields have not been extensively explored. In particular, the fields of language sciences, cognitive research, and social sciences could greatly benefit from targeted studies examining how machine learning and statistical analysis can uniquely advance our understanding of complex phenomena within these disciplines. Such investigations could elucidate the distinct advantages or limitations of each method, potentially leading to more refined and effective research methodologies in these areas.
The current study aims to bridge these gaps by employing both statistical analysis and machine learning on the same dataset to achieve different objectives. This study is intended to provide fresh perspectives to researchers in language sciences and cognitive studies, illustrating distinct applications of each method to extract unique insights.
2 Relation and differences between statistics and machine learning
Before starting our demonstration, we should systematically have a detailed account of how statistics and machine learning are related and differ from each other in theoretical perspectives. After having an understanding of these differences on theory, we believe that the practical implementation abilities will be enhanced, and this is our purpose of this paper.
@@ 2.1 How are statistics and machine learning related?
Breiman, 2001b highlighted the divergent approaches to data analysis, emphasizing the ongoing evolution of these methodologies over the decades. Many machine learning techniques have their roots in traditional statistical methods, such as linear and logistic regression, while also drawing from disciplines like calculus, linear algebra, and computer science. This intersection has led some to mistakenly merge the concepts of machine learning and statistics. Moreover, the introduction of user-friendly machine learning packages, such as Python’s scikit-learn (Pedregosa et al., , 2011), has further abstracted machine learning from its statistical foundation. This has propagated a belief among some newcomers to the field that a deep understanding of statistics isn’t necessary for machine learning applications. While basic tasks might not require intensive statistical knowledge, advanced modeling and the development of new algorithms heavily rely on a solid grounding in statistics and probability theory.
Statistical learning theory, formulated in the 1960s, lays the theoretical foundation for machine learning. It introduces concepts such as the hypothesis space and loss functions, providing a framework for supervised learning. The theory defines a dataset: Let $S$ be the set of all pairs $(x_{i},y_{i})$, where $x_{i}$ and $y_{i}$ are elements related by some function or condition. We then define $S$ as $S=\{(x_{i},y_{i})\}$, comprising $n$ data points, where each data point consists of features $x$ and a corresponding output $y$. The objective in statistical learning is to discover the function that maps the input features to the output, navigating through a hypothesis space of potential functions, guided by the minimization of a loss function which evaluates the expected risk over the dataset.
@@ 2.2 How are two fields different?
Despite the original intentions to remain separate, the core components of machine learning algorithms often rely on statistical principles that have been examined by statisticians for over a century. These mathematical principles apply universally, indifferent to whether the objective is to achieve artificial intelligence, publish research, or develop unbiased estimators. Consequently, many pressing questions in machine learning are, at their heart, statistical challenges previously unexplored by mainstream statistics. In recent decades, the field of statistics has found itself both challenged and invigorated by the successes of machine learning, particularly in areas traditionally dominated by statistics, such as predictive modeling (L’heureux et al., , 2017; Saidulu and Sasikala, , 2017; Ratner, , 2017; Rudin et al., , 2022; Ratner, , 2017; Ley et al., , 2022). This has sparked vigorous efforts to merge the theories and tools of both fields. However, for these efforts to be successful, they must first recognize and address the underlying reasons for their differences (Bzdok et al., , 2018; Makridakis et al., , 2018; Boulesteix and Schmid, , 2014). Understanding these differences is essential for leveraging the strengths of both fields effectively, whether the goal is to uncover deep insights from data or to develop robust predictive algorithms.
One of the fundamental distinctions between statistics and machine learning is their purpose. Statistics aims to infer properties about a population through samples, focusing on understanding and describing data relationships. In contrast, machine learning seeks to predict outcomes based on patterns identified in data, often using large and complex datasets to train predictive models. This training process typically involves dividing data into subsets for training, validation, and testing, which helps refine the models for better accuracy.
Another significant difference lies in how data is approached. In statistics, the focus is on the quality of data and the validity of the conclusions drawn from it through significance testing, which considers the presence of noise and potential confounding variables. Machine learning, however, emphasizes the quantity of data, often requiring large datasets to achieve the accuracy needed for effective predictions.
@@
2.3 An example
Linear regression is fundamentally a statistical method designed to minimize the squared error between data points. However, it is also widely used in machine learning for predictive tasks. An example of employing regression models is taken to illustrate different intentions and outcomes in language sciences.
In language sciences, regression models can be particularly illuminative. For example, researchers might use regression analysis to explore how the complexity of syntactic structures in children’s language development correlates with cognitive development indicators. Statistically, this would involve collecting data on children’s language use and cognitive tests, then applying linear regression to understand and quantify how changes in one variable relate to changes in another, typically across a complete dataset. The statistical approach focuses on inference—determining whether there is a statistically significant relationship between syntactic complexity and cognitive development. This method does not necessitate training and testing subsets of data; instead, it aims to characterize the relationship across the entire data set, assessing the significance and reliability of the observed relationships.
Conversely, in a machine learning context, the same regression model could be employed to predict cognitive developmental outcomes based on the existing linguistic data. The model would be trained on a subset of data, with the model’s parameters adjusted to optimize performance as measured on a separate test set. The objective is less about understanding the underlying dynamics of the relationship and more about achieving high predictive accuracy.
Suppose a language scientist wants to predict future cognitive development based on early linguistic behaviors using machine learning. In that case, they might employ a regression model on a divided dataset (training and testing), focusing on how well the model predicts developmental outcomes in new, unseen data. This approach would likely prioritize predictive power over interpretability, with model adjustments driven by performance metrics on the test set. Thus, while both fields use regression models, the context and objectives dictate their implementation: statistical modeling seeks to uncover and explain relationships within the data, providing comprehensive insights into language development patterns. In contrast, machine learning leverages these models primarily for their predictive capabilities, often in applications that require rapid assessments of developmental trajectories based on linguistic inputs.
These differences underscore the importance of distinguishing between the purposes and methodologies of statistics and machine learning in language sciences, ensuring that the chosen approach aligns with the specific goals of the research or application.
2.4 Key differences
In short, the contrast between machine learning and statistics is rooted in their different approaches and priorities. These key differences are summarized as shown in Table 1 and Fig. 1. The following provides some details on these key differences.
Figure 1: Differences between machine learning and statistics
Machine learning is primarily focused on achieving high predictive accuracy and is comfortable with models that are effective but not easily interpretable, often referred to as ‘black box’ models. On the other hand, statistics places a greater emphasis on relation among variables in data. Statisticians value models that are transparent and explainable, reflecting the field’s deep roots in mathematics and science, where theoretical foundations and provable properties are paramount. This includes a strong focus on the behavior of estimates as sample sizes increase and ensuring that models are robust even with small datasets.
Moreover, statistical methods rigorously address the bias and variance of estimates, seeking models that not only perform well but also provide insights into the certainty and reliability of predictions. In contrast, while machine learning does consider these factors, the emphasis is more on how the model performs in practical scenarios, often prioritizing larger datasets that feed the algorithmic complexity necessary for modern applications.
Table 1: Key Differences between Machine Learning and Statistics
@@ 2.4 Key differences
bias and variance of estimates Less Concerned Highly Concerned
@@ 3 Experiments
After having a systematic understanding the differences between statistics and machine learning, we carried out the experiments to demonstrate how machine learning and statistics differently contribute to linguistic and cognitive research.
@@ 3.1 Materials & methods
The Buckeye Corpus of conversational speech contains high-quality recordings from 40 speakers in Columbus OH conversing freely with an interviewer (Pitt et al., , 2005). The speech has been orthographically transcribed and phonetically labeled. The sessions were conducted as sociolinguistics interviews, and are essentially monologues. The speech has been orthographically transcribed and phonetically labeled. The corpus includes 357908 words.
The original dataset from the Buckeye provided diverse factors: speaker’s gender (female, male), age (old, young), word duration, etc. However, we can add more factors such as PoS tag for each word, word length, word frequency, the phrase rate, deletions and semantic relevance. Most of these factors have been investigate to show that they are closely related with word duration (Cowan et al., , 1997; Baker and Bradlow, , 2009; Terroir and Lavandier, , 2014; Cohen Priva, , 2015; Pierrehumbert, , 2016; Bürki, , 2018). However, these relevant research mostly employed statistical correlation and simply linear regression to explore how these factors affect word duration in speech. We adopt machine learning methods and some advanced mixed-effect regression models to examine them in the current study. The specific information on these factors is detailed in the Table 2:
Table 2: Factors and Their Descriptions
Factor Description
@@
Word Frequency (log) (Logarithm of) the normalized frequency of the word in the subtitle corpus  [^1] https://invokeit.wordpress.com/frequency-word-lists/.
CiteLength The number of syllables in transcript phonetic form of this word.
PhraseRate [Word number in this phrase] / [duration from the beginning of the phrase to the end of the phrase in a target word].
Deletion The number of segments in a target word deleted or reduced.
Semantic Relevance The semantic relatedness degree with the context.
@@
Additionally, semantic relevance is a novel measures, representing how a target word is semantically related with the context. It measures the semantic degree of how the contextual information influence the target word (Sun et al., , 2023; Sun, , 2023).
3.2 Machine learning methods

### v2

@@ Abstract
Data-driven approaches have revolutionized scientific research, with machine learning and statistical analysis being commonly used methodologies. Despite their widespread use, these approaches differ significantly in their techniques, objectives and implementations. Few studies have systematically applied both methods to identical datasets to highlight potential differences, particularly in language and cognitive sciences. This study employs the Buckeye Speech Corpus to illustrate how machine learning and statistical analysis are applied in data-driven research to obtain distinct insights on language production. We demonstrate the theoretical differences, implementation steps, and unique objectives of each approach through a comprehensive, tutorial-like comparison. Our analysis reveals that while machine learning excels at pattern recognition and prediction, statistical methods provide deeper insights into relationships between variables. The study highlights how semantic relevance, a novel metric measuring contextual influence on target words, contributes to understanding word duration in speech. We also systematically compare the differences between regression models used in machine learning and statistical analysis, particularly focusing on the training and fitting processes. Additionally, we clarify several common misconceptions that contribute to the confusion between these two approaches. Overall, by elucidating the complementary strengths of machine learning and statistics, this research enhances our understanding of diverse data-driven strategies in language and cognitive sciences, offering researchers valuable guidance on when and how to effectively apply these approaches in different research contexts.
Keywords: data-driven approaches, language production, objectives, prediction, features, implementations
@@ 1 Introduction
Advancements in technology have revolutionized the ability of scientists to gather and analyze data on an unprecedented scale. This surge in data availability has led to the necessity for more sophisticated quantitative techniques and computational tools, fundamentally changing the way research is conducted and facilitating new discoveries across various domains (Langley, , 1981; Han et al., , 1993; Kitchin, , 2014; Montáns et al., , 2019; Jack et al., , 2018; Núñez et al., , 2019). Data-based technologies serve as powerful illuminators, revealing hidden patterns and insights within vast datasets, much like a flashlight cutting through darkness. For example, machine learning (ML) has become indispensable in scientific inquiry, allowing researchers to decipher complex patterns, enhance their understanding of intricate phenomena, and make precise predictions. This technology is applied across diverse fields including medicine, astronomy, genomics, social sciences, and environmental studies, enabling researchers to swiftly extract significant insights from large volumes of data, much faster than traditional methods would permit (Mjolsness and DeCoste, , 2001; Jordan and Mitchell, , 2015; LeCun et al., , 2015; Webb et al., , 2018).
Such acceleration not only deepens the exploration of research questions but also unveils knowledge that was once concealed or unreachable. On the other hand, statistical methods play a critical role in the modern landscape of scientific research. They are integral in every phase of a study, from planning and design to data collection, analysis, and the interpretation and reporting of results (Carleo et al., , 2019; Núñez et al., , 2019; Butler et al., , 2018; Grimmer et al., , 2021). While machine learning helps in identifying patterns that may elude human detection in massive datasets, statistical analysis continues to enrich scientific research by providing robust frameworks for making inferences and validating findings (Fisher, , 1955; Box, , 1976; Nelder, , 1986; Nosek et al., , 2012). Together, these tools are taken as powerful data-science tools, transforming scientific paradigms, propelling forward our capacity to understand and manipulate the natural world.
@@
Data-driven research has made significant strides across various fields, bringing to the forefront disciplines such as data science, statistics, machine learning, and deep learning (Solomatine and Ostfeld, , 2008; Wolf, , 2010; Miller and Goodchild, , 2015; Zhang et al., , 2011). Among these, machine learning and statistics form the foundational core in data science. While often conflated due to their shared capability to process data, machine learning and statistics frequently overlap, creating ambiguity around their distinct roles and unique contributions. Moreover, both machine learning and statistics take advantage of the shared knowledge of probability theory, linear algebra etc. Contemporary research methodologies frequently incorporate machine learning and statistical analysis, often employing both to harness the strengths of each approach. Many researchers have found it challenging to distinguish clearly between these two areas. Although both fields are fundamentally concerned with extracting knowledge from data, they differ significantly in their objectives, approaches and implementations. Gaining a clear understanding of these differences is crucial for mastering both fields and effectively applying appropriate data-driven strategies.
@@
Despite the prevalent combined use of the two methods, there remains a significant gap in studies that systematically apply both techniques to identical datasets to highlight potential differences in outcomes. The systematic comparisons between them have been rarely taken in social sciences and behavior research, and such comparisons could provide valuable insights into how each method processes and interprets the same dataset differently. Moreover, the specific contributions of these techniques to these fields have not been extensively explored. In particular, the fields of language sciences, cognitive research, and social sciences could greatly benefit from targeted studies examining how machine learning and statistical analysis can uniquely advance our understanding of complex phenomena within these disciplines. Such investigations could elucidate the distinct advantages or limitations of each method, potentially leading to more refined and effective data-driven methodologies in these areas.
The present study aims to bridge existing gaps in research methodology by applying both statistical analysis and ML techniques to the same dataset. By demonstrating theoretical differences, distinct implementation steps and achieving different objectives, we provide a comprehensive, tutorial-like comparison of these approaches. This detailed examination offers researchers in cognitive and social sciences not only theoretical insights into the differences between machine learning and statistics but also hands-on experience in implementing these methods. The current study illustrates the unique applications of each method in addressing relevant questions, implementing comprehensive procedures, and extracting valuable insights. By doing so, we aim to enhance the understanding of when and how to effectively apply these data-driven approaches in various research contexts for both researchers and practitioners. This study not only elucidates the theoretical differences between the two data-driven approaches but also provides effective practical implementations for each. It will be particularly valuable for empirical and data-driven research in the social and cognitive sciences.
2 Background: Relation and differences between statistics and machine learning
Before delving into our demonstration, it is crucial to systematically examine the relationship between statistics and ML, highlighting both their connections and distinctions from a theoretical perspective. By thoroughly understanding these differences, we believe that the practical implementations will be significantly enhanced. This foundational knowledge will provide a solid framework for applying these concepts effectively in real-world scenario.
@@ 2.1 How are statistics and machine learning related?
The early 20th century saw statistics rapidly advance, empowering scientists to quantitatively assess hypotheses like “Does treatment X affect outcome Y?”. This led to the development of tools focused on hypothesis testing and confidence intervals, aimed at providing clear answers and estimating effect sizes. In contrast, machine learning emerged from engineering with the goal of enhancing machine capabilities. Initially, it focused on improving machine functionality and, secondarily, on understanding intelligence itself. Unlike statistics, machine learning was not primarily concerned with verifying real-world truths; instead, these were seen as potential components of intelligent behaviors–a perspective still debated today. This fundamental difference in origins and objectives highlights the distinct approaches of statistics and machine learning, despite their overlapping applications in data analysis and prediction.
Breiman, 2001b highlighted the divergent approaches to data analysis, emphasizing the ongoing evolution of these methodologies over the decades. Many machine learning techniques have their roots in traditional statistical methods, such as linear and logistic regression, while also drawing from disciplines like calculus, linear algebra, and computer science. This intersection has led some to mistakenly merge the concepts of machine learning and statistics. Even though they started out separately, machine learning often uses statistical ideas that have been around for over 100 years. These math principles work the same whether you’re trying to invent artificial intelligence, do research, or create fair ways to measure things. Because of this, many big questions in machine learning are actually old statistical problems that have not been looked at much before.
Moreover, the introduction of user-friendly machine learning packages, such as Python’s scikit-learn (Pedregosa et al., , 2011), has further abstracted machine learning from its statistical foundation. This has propagated a belief among some newcomers to the field that a deep understanding of statistics is not necessary for machine learning applications. While basic tasks might not require intensive statistical knowledge, advanced modeling and the development of new algorithms heavily rely on a solid grounding in statistics and probability theory.
@@ 2.2 How are two fields different?
In recent years, statistics has been both pushed and energized by machine learning’s success, especially in areas like predicting outcomes, which used to be mainly statistics’ territory (L’heureux et al., , 2017; Saidulu and Sasikala, , 2017; Ratner, , 2017; Rudin et al., , 2022; Ratner, , 2017; Ley et al., , 2022). This has sparked vigorous efforts to merge the theories and tools of both fields. However, for these efforts to be successful, they must first recognize and address the underlying reasons for their differences (Bzdok et al., , 2018; Makridakis et al., , 2018; Boulesteix and Schmid, , 2014). Understanding these differences is essential for employing the strengths of both fields effectively, whether the goal is to discover deep insights from data or to develop robust predictive algorithms.
One of the fundamental distinctions between statistics and machine learning is their purpose. Statistics aims to infer properties about a population through samples, focusing on understanding and describing the relationships among variables in data. In contrast, machine learning seeks to predict outcomes based on patterns identified in data, often using large and complex datasets to train predictive models. This training process typically involves dividing data into subsets for training, validation, and testing, which helps refine the models for better accuracy.
Another significant difference lies in how data is approached. In statistics, the focus is on the quality of data and the validity of the conclusions drawn from it through significance testing. Advanced statistical models need to learn from the data by fitting. In contrast, machine learning, however, emphasizes the quantity of data, often requiring large datasets to train models to achieve the accuracy needed for effective predictions. Despite this, the fitting and training processes differ in their objectives, strategies, and evaluation criteria (More details can be seen in Discussion Section 4.2).
They employ distinct evaluation standards and validation methods. Statistical regression commonly uses metrics like p-values, R-squared, AIC (Akaike Information Criterion), and confidence intervals. In contrast, machine learning often uses metrics like accuracy, precision, recall, F1-score for classification, or RMSE (Root Mean Square Error) for regression tasks. Statistical analysis often relies on assumptions and theoretical properties of the model. However, machine learning emphasizes empirical validation, using techniques like cross-validation and hold-out test sets.
@@
In short, the contrast between machine learning and statistics is rooted in their different objectives, approaches, evaluation standards, and priorities. The key differences are summarized as shown in (a) of Table 1 and Fig.1.
Figure 1: Differences between machine learning and statistics. Note: Fig. 1 could be aligned with Table 1, Table 3, and Table 4, making it easier to understand it.
The following provides further key differences on technique characteristics. Machine learning is primarily focused on achieving high predictive accuracy and is comfortable with models that are effective but not easily interpretable, often referred to as “black box” models. On the other hand, statistics places a greater emphasis on relation among variables in data. Statisticians value models that are transparent and explainable, reflecting the field’s deep roots in mathematics and science, where theoretical foundations and provable properties are paramount. This includes a strong focus on the behavior of estimates as sample sizes increase and ensuring that models are robust even with small datasets. This is summarized in (b) of Table 1.
Table 1: Key Differences between ML and Statistics
(a) Differences on Implementing ML and Statistics
Procedure Machine Learning Statistics
Purpose Patterns for classification, clustering and regression Relationships among variables
Approach Train ML models Fit in statistical models
Evaluation Precision, recall, F1-score p-value, AIC
Outcome ML model to automatically analyze new data Clarify the relationship among variables
(b) Technical Differences
@@
2.3 Exemplification: Children’s cognitive/language development
Linear regression is fundamentally a statistical method designed to minimize the squared error between data points. However, linear regression is also widely used in machine learning for predictive tasks. An example of employing regression models in machine learning or statistical analysis is taken to illustrate different intentions and outcomes in language sciences.
In language sciences, regression models can be particularly illuminative. For example, researchers might use regression analysis to explore how the complexity of syntactic structures in children’s language development (or other linguistic factors, such as vocabulary diversity, sentence length, speech rate or fluency measures) correlates with cognitive development indicators (indicated by scores on cognitive tests) (Goldin-Meadow et al., , 2014; Ünal and Papafragou, , 2020; Monshizadeh et al., , 2021). Statistically, this would involve collecting data on children’s language use and cognitive tests, then applying linear regression to understand and quantify how changes in complexity of syntactic structure relate to children cognitive development, typically across a complete dataset. The statistical approach focuses on inference—determining whether there is a statistically significant relationship between syntactic complexity and cognitive development. This method does not necessitate training and testing subsets of data; instead, it aims to characterize the relationship across the entire data set, assessing the significance and reliability of the observed relationships.
Conversely, in a machine learning context, the same regression model could be employed to predict cognitive developmental outcomes based on the existing linguistic data (e.g., syntactic complexity, age, vocabulary diversity, sentence length, speech rate or fluency measures etc.) (Briglia et al., , 2021). The model would be trained on a subset of data, with the model’s parameters adjusted to optimize performance as measured on a separate test set. The objective is less about understanding the underlying dynamics of the relationship and more about achieving high predictive accuracy.
Suppose a language scientist wants to predict future cognitive development based on early linguistic behaviors using machine learning. In that case, they might employ a regression model on a divided dataset (training and testing), focusing on how well the model predicts developmental outcomes in new, unseen data. This approach would likely prioritize predictive power over interpretability, with model adjustments driven by performance metrics on the test set. Thus, while both fields use regression models, the context and objectives dictate their implementation: statistical analysis seeks to uncover and explain relationships within the data, providing comprehensive insights into language development laws. In contrast, machine learning employs regression models primarily for their predictive capabilities, often in applications that require rapid assessments of developmental trends based on linguistic inputs.
These distinctions highlight the importance of differentiating between the aims and approaches of statistics and machine learning in exploring linguistic and cognitive questions. It is crucial to ensure that the chosen approach aligns with the specific goals of the research or application. In the following Experiment section, we will detail the implementation procedures, objectives, and outcomes of both machine learning and statistical analysis when applied to the same dataset. This comparison will provide a clearer understanding of how regression models are used differently in this example: studying children’s language development. In the Discussion section, we will further explore the distinctions between regression models employed in machine learning and statistical analysis, highlighting their unique characteristics and applications.
@@ 3 Experiments
Having systematically explored the theoretical differences between machine learning and statistics, we conducted experiments to demonstrate their distinct implementations and unique contributions to linguistic and cognitive research. These practical demonstrations aim to illustrate how each approach is implemented and the specific insights they offer in these fields.
@@ 3.1 Materials & methods
The Buckeye Corpus of conversational speech contains high-quality recordings from 40 speakers in Columbus OH conversing freely with an interviewer (Pitt et al., , 2005). The speech has been orthographically transcribed and phonetically labeled. The sessions were conducted as sociolinguistics interviews, and are essentially monologues. The speech has been orthographically transcribed and phonetically labeled. The corpus includes 357,908 words.
The original dataset from the Buckeye corpus provided diverse factors: speaker’s gender (female, male), age (old, young), word duration, etc. However, we can add more factors such as PoS tag for each word, word length, word frequency, the phrase rate, deletions and semantic relevance. Most of these factors have been investigated to show that they are closely related with word duration (Cowan et al., , 1997; Baker and Bradlow, , 2009; Terroir and Lavandier, , 2014; Cohen Priva, , 2015; Pierrehumbert, , 2016; Bürki, , 2018). However, these relevant research mostly employed statistical correlation and simply linear regression to explore how these factors affect word duration in speech. In this study, we employ both machine learning techniques and advanced mixed-effect regression models to analyze the data in the Buckeye corpus, aiming to gain complementary insights from these distinct approaches. The specific information on these factors is detailed in the Table 2:
Table 2: Factors and their descriptions in the the Buckeye corpus
Factor/Feature Description
@@
(log) Word Frequency (Logarithm of) the normalized frequency of the word in the subtitle corpus  [^1] https://invokeit.wordpress.com/frequency-word-lists/.
CiteLength (i.e., Syllable Number) The number of syllables in transcript phonetic form of this word.
PhraseRate (related to Speaking Rate) [Word number in this phrase] / [duration from the beginning of the phrase to the end of the phrase in a target word].
Deletion (related to Phonological Reduction) The number of segments in a target word deleted or reduced.
Semantic Relevance The semantic relatedness degree between the target word and its context (see the Appendix A).
@@
Additionally, semantic relevance is a novel metrics, representing how a target word is semantically related with the context. It measures the semantic degree of how the contextual information influence the target word (Sun et al., , 2023; Sun, , 2023). This metrics we introduced has proven effective in predicting eye-movements during reading multiple languages (Sun, , 2023), which encompasses a facet of language comprehension. This metrics is also capable of predicting and elucidating neural activity associated with the processing of naturalistic discourse, such as electroencephalography (EEG) (Sun et al., , 2024). The computation of semantic relevance is seen in the Appendix A.
3.2 Study 1: Machine learning methods